# A Time-Intensity Aware Pipeline for Generating Late-Stage Breast DCE-MRI using Generative Adversarial Models

### v1

@@ 4.6 Models Comparison
Finally, we evaluated the PAN model performance and compare it with the CE-PAN and TI-PAN using the scores described in section 4.2. Table 1 shows the comparative performance between all models in 1.5T and 3T images.
In the case of 1.5T images, pixels metrics are very similar, and best performances are divided for PAN (SSIM full image, PSNR and SSIM ROIs) and TI-PAN (MAE and PSNR full image). These results align correctly with the qualitative images shown previously. Similarly, in the case of the 3T images, CE-PAN obtains the best performance in most metrics, except PSNR in ROIs. Nevertheless, the difference margin between metrics is not large, and most models achieve similar values, where the largest difference is obtained in the PSNR with 0.3 and 0.1 for 3T and 1.5T respectively. However, the case of the TI metrics is different, since TI-PAN obtains the best scores compared to PAN and CE-PAN. In the case of annotated ROIs extracted from tissue, the TI-PAN increases the estimation of CE intensity by $0.05\%$ in the case of 1.5T and $0.03\%$ in the case of 3T. Although the increase is not large, it might be crucial in the diagnosis of patients since it might represent the difference between the different types of curves within the tissue.
A similar behavior is obtained for the unannotated regions, where an increase in CE intensity estimation is obtained for 1.5T images ($0.11\%$) and 3T images ($0.07\%$). This result is relevant because it shows the ability of our model to estimate the intensity of CE, regardless of whether an anomaly is present in the tissue or not. This also shows that the TI-PAN model does not rely on the presence of tumor, but leverages the biological behavior of the CA in tissue. Finally, a large margin is reported in the $E_{ps}$ score, where an increase of $0.20\%$ for both 1.5T and 3T images. The increase in performace is caused by the TI-Norm and TI-Loss altogether, since they exploit the TI curve to synthesize the images.
@@ 5 Discussion
Late enhancement, which is performed 8 or 10 min after injection of the contrast agent, provides relevant information to distinguish benign from probably malignant lesions in a Breast MRI study. However, acquiring it has the drawback of additional acquisition time and costs. Abbreviated protocols have gained attention as a cost-effective alternative in the medical field. However, concerns about their accuracy persist due to the omission of the late contrast enhancement phase, which, as mentioned, is important for an accurate diagnosis. Therefore, the development of alternatives for estimating the late contrast enhancement response is a valuable tool for integrating abbreviated protocols into clinical practice.
@@
In this work, we propose a pipeline that uses a deep learning-based generative model to predict the late response of a DCE-MRI from its early counterpart, which is optimized to preserve the expected biological behavior of the contrast agent by introducing the Time-Intensity loss function. The pipeline leverages the behavior of contrast agents at multiple levels, demonstrating its ability to reliably outperform the synthesis of clinically relevant information while preserving the visual properties of the images compared to existing models.
@@
To our knowledge, there are no previous works that have addressed the problem of the synthesis or generation of late-enhanced images for breast DCE-MRI. However, some works that attempt to generate contrast-enhanced images have been reported. Those works have shown adequate performance to generate realistic images with acceptable performances from a low-level feature point of view. However, no evaluation of the ability to reproduce the behavior of the contrast agent pattern has been considered, which is the key factor in determining the feasibility of using it in clinical practice. The results presented in this work demonstrate that image quality metrics are not completely suitable to optimize clinical information beyond visual similarity. Instead, task-specific types of functions, such as the proposed TI-Loss, could help achieve clinically interpretable results.
@@
The synthesizer performance was evaluated two ways: using stablished pixel-wise metrics, and evaluating the ability to replicate clinically relevant information (TI curve). Experimental results for the pixel-wise metrics show enough realism and similar performance in the full image setting, and also in annotated ROIs and unannotated regions. For the second group, we formulate two new metrics named the CA pattern score ($\mathcal{CP}_{s}$) and the average difference in the percentage of enhancement ($\mathcal{ED}$). In this case, the TI-PAN results significantly outperform other models by a large margin in both metrics. These results are relevant since they demonstrate that our pipeline is able to maintain image quality during the synthesis process while outperforming the clinical performance of the model and preserving the full interpretability of the diagnostic information.
@@
Image normalization is a standard stage in any image synthesis process. We evaluated multiple normalization techniques commonly used in the field of medical image synthesis and found that the most common normalization approaches (i.e. Min-max and Z-score) significantly modify the behavior of the TI curve, destroying its objective interpretation, as radiologists do in their daily practice. For this reason, we introduced the time-intensity normalization (TI-norm) strategy that preserves the diagnostic information required for physician interpretation. Our results in enhanced selected regions show that the proposed TI-norm preserves the information in the TI curve beyond the visual cues.
@@
This work differs from others because we proposed an entire pipeline where all stages were carefully intervened in to preserve important information for the clinical scenario, rather than only a single component, a model or function to process the data. Experimental results are promising for the implementation of late contrast-enhanced synthesis in a clinical setting. However, it must be evaluated by a controlled clinical study to determine the validity of the diagnostic information in a real assessment, particularly under different acquisition variables, variability among clinicians, inter-reader interpretation, etc. In particular, performance with different contrast agents and dosages must be evaluated, considering that the pipeline is grounded in the qualification of the CA response and the TI-curve. These variations cause differences in visualization and might alter the training process. This kind of study is out of the scope of this work and will be evaluated as a future line of research.
@@ 6 Conclusion
In this work, we introduced a strategy for predicting late-contrast-enhanced images of its early counterpart in breast DCE-MRI. We have proven the discrepancy between image quality during the synthesis and the prevalence of clinically significant information determined by contrast enhancement behavior. Based on this concept, we introduce three main innovations for clinical image synthesis: 1) a novel loss function that leverages the biological behavior of the contrast agent (CA) in tissue to optimize generative models, 2) a normalization strategy based on the time-intensity (TI) curve that preserves the diagnostic information required for interpretation and diagnosis, and 3) two performance evaluation metrics that measure how synthesized late enhanced images are able to replicate the TI-curve in enhanced regions. The proposed strategy demonstrates suitability for image synthesis by preserving the image quality of models but significantly outperforming the interpretability of the TI curve, which showed promising potential in clinical applications.
Funding
This work was supported by Minciencias, Colombia under the “programa de becas de excelencia doctoral del bicentenario”, BPIN 2021000100029.

### v2

@@ 3.6 Models Comparison
Finally, we evaluated the performance of the PAN model and compared it with the CE-PAN and TI-PAN using the scores described in section 3.2. Table 1 shows the comparative performance among all models in 1.5T and 3T images.
In the case of 1.5T images, pixel metrics are very similar, and the best performances are divided between PAN (SSIM full image, PSNR and SSIM ROIs) and TI-PAN (MAE and PSNR full image). These results align well with the qualitative images shown previously. Similarly, in the case of the 3T images, CE-PAN obtains the best performance in most metrics, except PSNR in ROIs. Nevertheless, the margin of difference between metrics is not large, and most models achieve similar values, with the largest difference obtained in PSNR, at 0.3 and 0.1 for 3T and 1.5T respectively. However, in the case of the TI metrics, significant differences were observed since TI-PAN obtains the best scores compared to PAN and CE-PAN. In the case of annotated ROIs extracted from tissue, the TI-PAN increases the estimation of CE intensity by $0.05\%$ in the case of 1.5T and $0.03\%$ in the case of 3T. Although the increase is not large, it might be crucial in the diagnosis of patients, as it might represent the difference between the different types of curves within the tissue.
A similar behavior is observed for the unannotated regions, where an increase in CE intensity estimation is noted for 1.5T images ($0.11\%$) and 3T images ($0.07\%$). This result is relevant because it demonstrates the ability of our model to estimate the intensity of CE, regardless of whether an anomaly is present in the tissue or not. This also indicates that the TI-PAN model does not rely on the presence of a tumor but rather leverages the biological behavior of the CA in the tissue. Finally, a large margin is reported in the $E_{ps}$ score, where an increase of $0.20\%$ is observed for both 1.5T and 3T images. The increase in performance is attributed to the TI-Norm and TI-Loss together, since they exploit the TI curve to synthesize the images.
@@ 4 Discussion
In breast DCE-MRI, late enhancement images, which can be acquired 8 or 10 minutes after the injection of the contrast agent, provide relevant information to distinguish benign from probably malignant lesions  [5]. However, acquiring these images has the drawback of additional acquisition time and costs. Although alternatives such as abbreviated protocols have gained attention in clinical scenarios, concerns about their accuracy persist due to the omission of the late contrast enhancement phase, critical for an accurate diagnosis  [37, 38, 39]. Therefore, the development of alternatives for estimating the late contrast enhancement response is a valuable tool for moving towards their applicability in clinical practice.
@@
In this work, we propose a pipeline that uses a deep learning-based generative model to predict the late response of a DCE-MRI from its early counterpart, which is optimized to replicate the biological behavior of the contrast agent by introducing the Time-Intensity loss function. The pipeline leverages the behavior of contrast agents at multiple levels, demonstrating its ability to reliably outperform the synthesis of clinically relevant information while preserving the visual properties of the images when compared to existing models.
@@
To our knowledge, there are no previous works that have addressed the problem of synthesis or generation of late-enhanced images for breast DCE-MRI. However, some works that attempt to generate contrast-enhanced images have been reported. Those works have demonstrated adequate performance in generating realistic images with acceptable quality from a low-level feature perspective. Nonetheless, no evaluation of the ability to reproduce the behavior of the contrast agent pattern has been considered, which is a key factor in determining the feasibility of using it in clinical practice. The results presented in this work demonstrate that image quality metrics are not completely suitable for optimizing clinical information beyond visual similarity. Instead, task-specific types of functions, such as the proposed TI-Loss, could help achieve clinically interpretable results.
@@
The performance of the synthesis was evaluated in two ways: using established pixel-wise metrics and assessing the ability to replicate clinically relevant information (TI curve). From the pixel-wise metrics perspective, experimental results showed sufficient realism and comparable performance in both the full image setting and selected regions (radiological findings and unannotated regions). On the other hand, from the clinical measures point of view, the proposed CA pattern ($\mathcal{CP}_{s}$) and the average enhancement differences ($\mathcal{ED}$) scores revealed significant differences that could seriously affect diagnostic decisions. In this case, the TI-PAN significantly outperformed other models, demonstrating that the proposed pipeline is able to maintain image quality during the synthesis process while replicating accurately the clinical performance of the model, preserving the full interpretability of the diagnostic information.
@@
Image normalization is a standard stage in any image synthesis process. We evaluated multiple normalization techniques commonly used in the field of medical image synthesis and found that the most common normalization approaches (i.e. Min-max and Z-score) significantly modify the behavior of the TI curve, distorting its objective interpretation. For this reason, we introduced the time-intensity normalization (TI-norm) strategy that preserves the diagnostic information required for physician interpretation. Our results in enhanced selected regions showed how the proposed TI-norm preserves the information in the TI curve beyond the visual cues.
@@
This work differs from others because we propose an entire pipeline in which all stages are carefully intervened to preserve important information for the clinical scenario, rather than focusing solely on a single component, model, or function to process the data. Experimental results are promising for the implementation of late contrast-enhanced synthesis in a clinical setting. However, it must be evaluated through a controlled clinical study to determine the validity of the diagnostic information in a real assessment, particularly under varying acquisition parameters, variability among clinicians, and inter-reader interpretation, etc. In particular, performance with different contrast agents and dosages must be evaluated, considering that the pipeline is grounded in the qualification of the CA response and the TI-curve. These variations can cause differences in visualization and may alter the training process. This type of study is beyond the scope of this work and will be considered as a future line of research.
@@ 5 Conclusion
This work presents a comprehensive pipeline for synthesizing late-phase contrast-enhanced breast MR images from the early-phase counterpart. Our approach leverages the temporal behavior of contrast agents through the introduction of the Time Intensity Loss (TI-loss) function and a novel normalization strategy (TI-norm), which together guide the training of a generative model to accurately replicate clinically relevant information while preserving spatial features across the entire image, outperforming models optimized solely for pixel-wise quality. The synthesized images maintained diagnostic utility comparable to real late-phase images, as evidenced by the proposed evaluation metrics. Experimental results demonstrated that established pixel-based metrics do not guarantee the prevalence of contrast agent behavior, which is critical in clinical applications.
By reducing the need for late-phase acquisitions, this pipeline has the potential to shorten scanning times, improve patient comfort, and alleviate cost and availability constraints associated with conventional breast DCE-MRI protocols. This advancement brings generative models closer to practical implementation in clinical scenarios, enhancing efficiency in breast cancer imaging without compromising diagnostic accuracy.
Future work must perform clinical validation through collaboration with healthcare professionals, which will be essential to ensure the method’s efficacy and safety in real-world settings.
Ethics statement
This study exclusively uses publicly available datasets, which are fully anonymized and do not contain identifiable personal or sensitive information. Datasets are accessible to the public and do not involve any interaction with human subjects.
Acknowledgements
This work was supported by Minciencias, Colombia under the “programa de becas de excelencia doctoral del bicentenario”, BPIN 2021000100029 and the Instituto Tecnológico Metropolitano.